OpenStudy (anonymous):

A particle executes SHM with frequency 2.5 Hz and amplitude 2 m. The speed of the particle 0.3 s after crossing, the equilibrium position is????

OpenStudy (anonymous):

please guys.... Help....

OpenStudy (anonymous):

y = (2) sin (5 pi t) because f = 2.5/s = 2.5 Hz freqeuency y' = dy/dt = (2)(5 pi) cos (5 pi t) velocity from derivative "equilibrium position" means? probably y =0 , so t=0 would work thus 0.3s = t is where to evaluate y' from above.

OpenStudy (anonymous):

Yesss....... Thanks..... I totally forgot about that....